OpenStudy (anonymous):

A certain box of cereal should have a net weight that is within 1.5 grams of 510. After the first phase of production, the weight of the cereal is 308 grams. By how much should the weight of the cereal increase to remain within the specified weight?

OpenStudy (aaronq):

510-308=

OpenStudy (anonymous):

the answer is 202?

OpenStudy (aaronq):

you're saying that the cereal box should weigh \[510 \pm 1.5 g \] correct? if you already have 308 g because of the weight of the box, how much weight should you add? 510-308= \[510-308=202\pm 1.5 g\]